OpenStudy (anonymous):

how do you graph y = x + 4

OpenStudy (anonymous):

its supposed to say y=-3/2x+4

OpenStudy (hba):

You can consider values of x and get values of y You can take upto 4 values of (x,y) and the plot it For example. x=1 y=1+4 y=5 (x,y)=(1,5)

OpenStudy (hba):

It would be the same for that too.

OpenStudy (hba):

Just consider values and get 4 values of (x,y) and plot them :)

OpenStudy (anonymous):

okayy thanks (: